A fund has $878 million dollars of assets, $21 million of liabilities, and 13 million shares outstanding. During the year the fund made qualified dividend distributions of $2.9 per share, long term capital gains distributions of $1.6, and short-term capital gains distributions of $1.9. What is the after-tax distribution yield, if your marginal income tax rate is 35%, and your long term capital gains tax rate (which also applies to qualified dividends) is 15%?

Enter answer in percents, accurate to two decimal places.

Extra: How does that compare with pre-tax yield? What are the criteria for dividends to be "qualified" under current tax laws?
Answer #1

solution NAV = = = ( assets - liabilities) (no. of shares outstanding ( 878 000 000 - 20000000) / 13000000 65.92 To compute pPretax earnings = 2.9 + 1.6 + 1.9 = 6.40 pretax field = pretax earnings / NAV = = 6.40 / 65.92 9.70 % 9. 70 % post tay earnin